OpenStudy (anonymous):

Solve for x: 5x + 2 = 4x − 9

OpenStudy (anonymous):

5x + 2 = 4x - 9 - 2 -2 5x= 4x - 11 -4x -4x 1x = -11

OpenStudy (anonymous):

Thank you Ethompsonn :}